# ClipRover: Zero-shot Vision-Language Exploration and Target Discovery by Mobile Robots

### v1

@@ 4.1 Visual Perception Frontend
In the frontend, raw camera frames are sliced into six; the slicing strategies are discussed in Sec. 5.1. Each tile represents a spatial location in robot’s FOV, which are scaled and processed by the CLIP’s visual encoder. As shown in Fig. 5 f, each frame is sliced into $N=6$ tiles and then rearranged into a tensor of shape $N\times 3\times H\times W$, with $H$ and $W$ are the tile height and width in pixels. The encoder processes these inputs and generates an $N\times D$ embedding vector for each tile, where $D$ is the dimensionality of each prediction vector ($D=512$ in the CLIP model).
@@
Figure 5: An overview of ClipRover architecture; the primary data flow and supplementary sensory feedback are in black and red arrows, respectively. The camera frame-slicing strategy is shown in (f); each frame is sliced into two rows (NEAR, FAR) and three columns (LEFT, CENTER, and RIGHT). Each of the six tiles is encoded into a separate visual embedding vector for perception.
@@ 4.3 Familiarity Middleware: Vision-Vision Correlation
When no known vector exists in the database, the incoming vector is inserted as a new data point. Eventually, a familiarity score is generated for each perception vector in the database. This score guides navigation by encouraging the robot to prioritize unexplored areas over revisiting familiar ones.
@@ 4.4 Navigation Decision Backend
4.4.1 Trap Detection
In certain scenarios, the robot may encounter a “dead end”, where no navigable path is visible within its FOV. In rare instances, the vision-language model may generate false positive Nav scores for scenes it does not adequately comprehend. For example, the lack of salient features can lead to incorrect positive scores when the camera is positioned too close to a plain white wall. Such situations are defined as “trapped” states, which can occur under two conditions: (a) the proximity switch asserts a halt signal for a specified duration, or (b) the cumulative travel distance, as measured by odometry, falls below a defined threshold over a given period. Empirically, the system flags the robot as trapped if it travels less than $0.2$ meters within the past $5$ seconds.
4.4.2 Look Around
Due to the limited FOV from a single camera, the robot could only see objects in front of it. To address this, a “look around” mechanism (see Fig. 8) is introduced to enable the robot to gain situational awareness by performing a $360^{\circ}$ rotation while collecting Nav scores associated with different headings. A Gaussian convolution is then applied to these scores to identify the most navigable direction. When recovering from a “trapped” state, the look around mechanism prioritizes a direction different from the original heading by a linear factor $k$, rewarding headings that deviate from the initial orientation. Additionally, a look around behavior is triggered at the start of a new mission to ensure the robot identifies the most promising path for exploration.
@@ 4.4.2 Look Around
5 Implementation Details
5.1 Slicing Strategies
@@
To enhance the embedding of positional information within the pipeline, we preprocess the raw camera frames by slicing them into smaller tiles before inputting them into the vision-language model. This slicing strategy is specifically designed to optimize the robot’s ability to make informed navigation decisions based on its visual perceptions. As illustrated in Fig. 1, the slicer divides each frame into six ($2~{}\text{rows}\times 3~{}\text{columns}$) tiles, with the number of horizontal divisions aligned with the robot’s control capabilities. Using smaller tiles instead of the full camera frame reduces visual distractions, thereby improving the accuracy of the correlation results.
@@
5.2 Prompt Generation
@@
We utilize a YAML-based syntax to define the structure of the prompt databases. YAML, a widely adopted and human-readable markup language, has been extended in our implementation to incorporate additional syntactical features for enhanced flexibility and fine-grained control over the generation of text prompt combinations. Two key features introduced are: Selective integration and in-place expansion.
@@
    In-place expansion facilitates the collective definition of similar short prompts by using a vertical bar syntax (i.e. |) to separate terms. This feature is particularly effective when used in conjunction with Selective Integration. For example, the prompt “A photo with no {context|texture|information}” is expanded into three distinct prompts, although not fitting into the hierarchical template structure, can be defined as top-level negative prompts, aiding in distinguishing meaningful content from irrelevant or meaningless information.

### v2

@@ 4.1 Visual Perception Frontend
In the frontend, raw camera frames are sliced into six tiles; the slicing strategies are discussed in Sec.1 of the supplementary materials. Each tile represents a spatial location in the robot’s FOV, which are scaled and processed by CLIP’s visual encoder. As shown in Fig. 5 f, each frame is sliced into $N=6$ tiles and then rearranged into a tensor of shape $N\times 3\times H\times W$, with $H$ and $W$ being the tile height and width in pixels. The encoder processes these inputs and generates an $N\times D$ embedding vector for each tile, where $D$ is the dimensionality of each prediction vector ($D=512$ in the CLIP model).
@@
Figure 5: An overview of the architecture of VL-Explore; the primary data flow and supplementary sensory feedback are marked by black and red arrows, respectively. The camera frame-slicing strategy is shown in (f); each frame is sliced into two rows (NEAR, FAR) and three columns (LEFT, CENTER, and RIGHT). Each of the six tiles is encoded into a separate visual embedding vector for perception.
@@ 4.2 Navigability Middleware: Vision-Language Correlation
The symbols used in examples above belong to a custom designed prompt system. The same set of notations are used throughout this paper. The symbols are explained in detail in Sec.2 of the supplementary materials.
@@ 4.3 Familiarity Middleware: Vision-Vision Correlation
When no known vector exists in the database, the incoming vector is inserted as a new data point. Eventually, a familiarity score is generated for each perception vector, guiding navigation by encouraging the robot to prioritize unexplored areas over revisiting familiar ones.
@@ 4.4 Navigation Decision Backend
Trap Detection. In certain scenarios, the robot may encounter a “dead end”, where no navigable path is visible within its FOV. In rare instances, the vision-language model may generate false positive Nav scores for scenes it does not adequately comprehend. For example, the lack of salient features can lead to incorrect positive scores when the camera is positioned too close to a plain white wall. Such situations are defined as “trapped” states, which can occur under two conditions: (a) the proximity switch asserts a halt signal for a specified duration, or (b) the cumulative travel distance, as measured by odometry, falls below a defined threshold over a given period. Empirically, the system flags the robot as trapped if it travels less than $0.2$ meters within the past $5$ seconds.
Look Around. Due to the limited FOV from a single camera, the robot could only see objects in front of it, limiting the amount of information available to the robot for efficient navigation and exploration. To address this, a “look around” mechanism (see Fig. 8) is introduced to enable the robot to gain situational awareness by performing a $360^{\circ}$ rotation while collecting Nav scores associated with different headings. A Gaussian convolution is then applied to these scores to identify the most navigable direction. When recovering from a “trapped” state, the look around mechanism prioritizes a direction different from the original heading by a linear factor $k$, rewarding headings that deviate from the initial orientation. Additionally, a look around behavior is triggered at the start of a new mission to ensure the robot identifies the most promising path for exploration.
@@
5 Experimental Results and Analyses
5.1 Experimental Setup
Real-world experiments were conducted in two different environments as depicted in Fig. 10 a. The environments were selected for their cluttered layout and complex details, which include numerous obstacles, potential traps, and loops, providing a challenging setting for comprehensive analysis. As shown in Fig. 10, the robot was tasked with exploring the space while searching for a designated target – a toy bear approximately $20$ cm tall and $10$ cm wide, chosen for its distinctive appearance within the test scene. For each task, the source (robot’s starting position) and destination (target location) were selected from predefined locations as labeled in Fig. 10 (b) and (d).
Figure 10: The environment used for real world exploration and discovery experiments: (a) Overview of the machine hall taken from the east side. (b) 2D map of the experimental space, with task locations annotated in blue; the area utilized for the experiments is enclosed within the dashed red line. (c) An office scene taken from the S. E.corner; (d) 2D map of the office scene, with task locations annotated in blue.
5.2 Performance Evaluation
Algorithms for comparison. We compare the performance of the proposed VL-Explore system with six widely used map-traversal and path-finding algorithms. For fair comparisons, a novel 2D simulation framework, RoboSim2D, is developed. It utilizes 2D LiDAR maps from recorded scans of the same environment used in real-world experiments. Additionally, the size of the simulated robot was configured to match the dimensions of the physical robot, ensuring consistency across simulations and the real platform. See the supplementary materials (Sec. 5) for more details on the algorithmic implementation and simulation results.
Criteria of failure. Due to practical limitations, a mission is considered a failure if the distance traveled exceeds a predefined upper limit. For real-world experiments conducted in the test area (approximately $12\times 16$ m), this limit is set to $100$ meters. In simulations, the limit is set to $1,000$ meters for random walk, wall-bouncing, and wave-front algorithms. For Bug algorithms, loop detection is employed to identify endless loops, which are classified as failures. Since failed missions yield infinite travel distances, they are excluded from the metrics presented in Table 2, they are instead reflected in the overall success rate of an algorithm.
$\text{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{l_{i}}{\max(p_{i},l_{i})}.$ (1)
Note that, for each source-target combination, the first contact distance of WaveFront simulation is used as the baseline distance $D_{\text{baseline}}$. With the choice of $l=D_{\text{baseline}}$, we have $p_{i}>l,\forall p_{i}\in\{p\}$; hence, the above equation is simplified to:
[…]
As a consequence, the SPL metric equates to an inverse proportional relation between two quantities, namely the success rate $R$ and the mean inverse path length $\overline{L}$, as:
$\text{SPL}=\overline{L}\cdot{R}.$ (4)
However, experimental data reveal that the curvature of the SPL function does not align well with the observed $R$$\overline{L}$ curve, shown later in the experimental results (Fig. 11). That is, for the same exploration algorithm, setting a different failure criterion will result in a different SPL score. To better model the efficiency of an algorithm from observed data, we propose a new metric “Entropy Preserving Score” (EPS) to represent the amount of information available to a given algorithm. It has a range of $[0,1]$. When $\mathbf{\text{EPS}}=0$, the algorithm operates without environmental information and lacks sensing capabilities beyond collision detection, as seen in methods like random walk and wall bounce. As EPS approaches $1$, it gains more information or has higher environment sensing capability, and the performance is expected to increase accordingly. Considering the impact of EPS, the revised equation expands to:
$f_{1}(\mathbf{\text{EPS}})=f_{2}(\overline{L})\cdot f_{3}(R)$ (5)
[…]
where $H_{n}=[k_{n}~{}~{}t_{n}~{}~{}p_{n}]$ are hyper-parameters specific to an environment. i.e. A specific environment can be characterized using the collection of 9 hyperparameters:
$H=\left[\begin{matrix}H_{1}\\ H_{2}\\ H_{3}\end{matrix}\right]=\left[\begin{matrix}k_{1}&t_{1}&p_{1}\\ k_{2}&t_{2}&p_{2}\\ k_{3}&t_{3}&p_{3}\end{matrix}\right].$ (6)
These hyper-parameters are fitted to the $\overline{L}-{R}$ curve derived from random-walk simulation results in the given map, plus an additional boundary condition $f_{1}(1.0)=f_{2}(1.0)\cdot f_{3}(1.0)$, which is the ideal case when the algorithm has all information available and achieves a perfect success rate with optimal travel distance.
Table 2: Quantitative performance of VL-Explore and other algorithms in comparison. Metrics used are defined in Eq. 3.
Table 4.  The acronyms: S.R. - Success Rate $R$; P.L. - Mean Inverse Path Length $\overline{L}$.
Figure 11: Performance of VL-Explore compared to map-traversal methods (i.e.  random walk, wall bounce) [34], path-finding (Bug) algorithms [28, 29] and SOTA VLN methods (i.e.  VLMnav and GOAT) [13, 6] for UGV navigation. SOTA VLN methods are represented by a solid gray dot (all systems enabled) and a gray circle (certain systems disabled to match with information required by VL-Explore). Detailed explanations are provided in Table.2. (a) Large industrial scene; (b) Mid-size office scene. The equipotential lines labeled as $\mathbf{\text{EPS}}=x$ are defined by Eq. 5.
Figure 12: Sample results of autonomous exploration and target discovery by VL-Explore compared to Bug algorithms. The trajectories of each algorithm are overlaid on a 2D map of the test environment (north up). Circled numbers indicate the look around operations of VL-Explore and the corresponding waypoints sequence it followed. For Bug algorithms, the red, green, and blue trajectories represent paths traversed by Bug0, Bug1, and Bug2, respectively. Best viewed digitally at 2×2$\times$2 × zoom for clarity; more results and analyses are included in Sec.6 of the supplementary materials.
5.3 Qualitative and Quantitative Analyses
We conduct extensive real-world experiments with over $60$ trials based on various combinations of source-target locations shown in Fig. 10. For the proposed VL-Explore system, three trials were performed for each origin-target pair. In comparison, $200$ trials were recorded for each random walk experiment, while $180$ evenly distributed headings were recorded for each combination of the wall bounce runs.
Samples of the trajectories traversed by VL-Explore and other algorithms are presented in Fig. 12, overlaid on a 2D map of the environment. Under normal navigation mode, VL-Explore exhibits a human-like motion strategy by navigating near the center of open spaces, enhancing exploration efficiency and reducing the likelihood of becoming trapped by obstacles. In contrast, path-finding algorithms often follow the contours of obstacles due to their lack of semantic scene understanding. During look-around operations (see Fig. 8 and Fig. 9), VL-Explore accurately discriminates between navigable spaces and obstacles. Additionally, it demonstrates a preference for unexplored (unfamiliar) areas over previously visited (familiar) regions, contributing to improved efficiency compared to traditional map-traversal algorithms.
Out of $60$ trials in the industrial scene, 3 failure cases were recorded, yielding a $95\%$ overall success rate for VL-Explore. For the office scene, only 1 failure case was recorded out of $12$ trials, yielding a $91.7\%$ success rate. Among the failure cases, two were caused by the trajectory length exceeding the limit, while the other two were caused by the robot getting jammed by an obstacle, which was not detected by either the VLN pipeline or the LiDAR proximity switch.
The quantitative results categorized by source and target locations are presented in Table.1 of the supplementary materials. To ensure comparability across different source-target pairs, the travel distance for each row is normalized by the baseline travel distance and then aggregated into Table 2. For randomized algorithms, the success rate $R$ is a configurable hyperparameter. Sample results for $R=50\%$ and $R=80\%$ are provided to represent their underlying performance in Table 2, with the corresponding equipotential curves previously shown in Fig. 11. In contrast, deterministic algorithms (such as Bug variants) have a fixed success rate. Their performances are reported directly in Table 2 and visualized in Fig. 11 for comparison.
As these results demonstrate, with the same amount of information, the proposed system outperforms traditional map traversal algorithms by significant margins in success rates, trajectory lengths, SPL and EPS scores. Despite the inherent disadvantage of operating without prior target knowledge or precise localization, the system achieves performance comparable to path-finding algorithms. Notably, in many scenarios, it surpasses path-finding algorithms in either trajectory efficiency or success rates. These results highlight the effectiveness of the proposed vision-language based pipeline for simultaneous exploration and target discovery tasks.
5.4 Comparative Analyses of SOTA VLNs
We also compare VL-Explore with SOTA VLN systems based on their features and prerequisites. It is important to note that, due to different experimental setup and reporting criteria, a perfect quantitative comparison is not always feasible. For example, SEEK [12] reported $0.65$±$\pm 0.35$ SPL for random walk, while our random walk simulation yielded $0.08$±$\pm 0.11$ SPL, showing a magnitude of difference. Preliminary analysis suggests that this discrepancy stems from SEEK’s reliance on a node-graph map structure, which reduces the problem space to a finite set of graph nodes and edges. In contrast, VL-Explore operates continuously in an unknown environment, resulting in a substantially larger problem space, facilitating active robot navigation with no prior information or pre-compiled waypoints. This fundamental difference motivated us to introduce the EPS metric, which helps reduce the discrepancies caused by varying prerequisites and experimental setups.
Among the vast volume of prior attempts in the field of vision-language based navigation or exploration, we select two recent representative systems that provide ablation results with similar prerequisites as our proposed system:
  • VLMnav [13] is an end-to-end VLN system, which assumes no prior knowledge or map; it does not rely on online SLAM for navigation. However, it requires a depth camera to obtain navigability scores (floor mask). It reported $0.210$ SPL on the ObjectNav dataset, its score drops to $0.136$ without the depth camera.
  • GOAT [6] is a lifelong mapping system; it requires an RGBD camera to obtain necessary information for navigation and map construction. In addition to the final target, it also takes “sequential goals” that help it to take intermediate moves. It reported $0.64$ SPL on the “in the wild” dataset with persistent memory, but the score drops to $0.19$ in the absence of prior knowledge. The authors did not perform an ablation study when depth sensing is removed.
In contrast, our proposed work, VL-Explore, achieves $\mathbf{0.357}$ SPL without requiring depth sensing, persistent memory, or prior knowledge (e.g.  mapping run) of a scene. To the best of our knowledge, VL-Explore is the first VLN system to achieve efficient indoor visual exploration without any depth or range information (e.g. RGBD, multi-camera, LiDAR). VL-Explore holds a significant advantage in its intended use case – first-time exploration and target discovery with a single monocular camera.
There also exist many other works based on different prerequisites, which cannot be directly included in quantitative comparison. We instead summarize the characteristics of each system in Table 3. Compared to existing systems, VL-Explore uniquely requires zero prior knowledge of the task space and operates solely with monocular RGB perception. Additionally, contemporary VLN systems do not directly produce low-level motion commands but instead generate high-level instructions that depend on another traditional localization and navigation pipeline for the eventual motion execution. In contrast, VL-Explore directly outputs low-level motion commands, eliminating the need for auxiliary control and localization systems. This feature makes VL-Explore particularly suitable for rapid deployment on low-cost robots with limited computational and sensing resources.
5.5 Ablation Study
Figure 13: Ablation study of the proposed VL-Explore  system; (a) Comparative results with all systems enabled; (b) Results with the look-around subsystem disabled, the robot either completes the task with no turn-around, or directly fail when being trapped, yielding a higher PL but a lower SR; (c) Results with the familiarity subsystem disabled, the robot tends to revisit previously explored areas, resulting in longer trajectories and lower success rates.
Table 3: A qualitative comparison of features and prerequisites among SOTA VLN systems, highlighting the unique advantage of the proposed VL-Explore  system that requires no extra sensor modalities or any separate path-finding or localization steps.
CLIP on Wheels [10] Object Unspecified RGB-D Camera ${}^{\text{Relevance map of}}_{\text{the current View}}$ Simulation ${}^{\text{Accuracy \&}}_{\text{Path length}}$
[…]
In order to understand the contribution of each middleware on the overall performance, we conduct an ablation study by disabling one middleware at a time and re-evaluating the system’s performance. The experiment is performed in the medium-sized office scene shown in Fig. 10 (c), with the robot starting from the South-West (SW) corner and tasked to find the target in the North-East (NE) corner.
  • Disabling the look-around action. In this part, the system does not perform any look-around action during a task, including the initial look-around that helps the robot finding its optimal initial heading. The initial orientation of the robot is manually set to a different direction for each trial. The results are shown in Fig. 13b. With the lack of look-around action, the task would fail upon encountering a dead-end as it could not recover from such a situation. This resulted in a significantly lower success rate ($50\%$). However, in the cases when the robot successfully navigated to the target, the reported inverse path lengths turned out better due to survivorship bias. i.e. inefficient runs tend to fail instead of being recovered by a look-around action.
  • Disabling the familiarity middleware. In this part, the familiarity middleware is disabled by truncating the familiarity scores to a fixed value. The results are shown in Fig. 13c. Without the familiarity middleware, the robot exhibited a strong tendency to revisit previously explored areas. In this particular case, the robot tends to run back and forth between S.W. and N.W. corner of the scene. This lead to significantly longer trajectory lengths (i.e. smaller mean inverse path length), and also induced negative impact on the overall success rate.
To summarize, the additional system components added to the VL-Explore pipeline are proven to be essential for its overall performance. In addition to the basic obstacle avoidance and target identification capabilities provided by the vision-language correlation middlewares, the look-around action helps the robot to optimize its decision and recover from trap conditions, and the familiarity middleware ensures efficient exploration by constantly steering the robot into unexplored area.
6 Limitations, Failure Cases, and Potential Improvements of VL-Explore
6.1 Texture-less or Transparent Surfaces
As a vision-driven system, the proposed VL-Explore  pipeline encounters challenges in environments with texture-less or transparent surfaces. When presented with a blank or non-informative image, the inherent VLM struggles to generate meaningful responses, disrupting the entire pipeline. For instance, as demonstrated in Fig. 9 d (C1), the robot mistakenly classified a water tank with transparent walls as navigable space. This error occurred because the VLM perceived the interior of the tank, visible through its transparent walls, as accessible terrain. To mitigate this issue, a standard deviation threshold was introduced to filter out ambiguous tiles. While this approach improved performance, the system still exhibits a general tendency to misclassify transparent or uniformly colored surfaces, such as blank walls or white floors, as navigable space. These challenges persist, particularly when the robot is positioned close to such surfaces, where the lack of texture further confuses the VLM. Note that this issue with transparent and texture-less walls has been an open problem for 2D indoor navigation by mobile robots [44].
6.2 Open Space and Artificially Constructed Mazes
Moreover, our experiments in artificially constructed mazes, such as those built from paper boxes, revealed limitations in the VLM’s ability to differentiate between the maze walls, such as monochrome cardboards and floors. As a result, the robot either failed to navigate or attempted to collide with the maze walls. This behavior can be attributed to the lack of similar examples in the VLM’s training data, which likely did not include artificially constructed environments of this nature. Consequently, the system is better suited to real-world scenarios where the VLM has been trained on relevant, diverse visual data. This observation suggests that our proposed system will work effectively in real-world environments where VLMs generally work well.
6.3 Familiarity Saturation
6.4 Adaptive Slicing Strategies
In our current implementation of VL-Explore, a fixed slicing strategy is used based on the camera frame’s aspect ratio, which generally performs well in general. However, in specific scenarios involving narrow navigable spaces (e.g., a $50$ cm gap between two obstacles), the robot often fails to recognize these spaces as traversable. This is because the narrow passage does not occupy an entire tile, leaving obstacles visible in all tiles and leading the system to avoid the space rather than navigate through it. Future improvements could include the development of advanced slicing strategies that dynamically adapt the region of interest based on the surroundings. This could be achieved using sparse object detection or pixel-wise segmentation models, enabling better handling of narrow spaces.
6.5 High Level Reasoning & Task Narration
One of the most promising extensions of this work lies in incorporating high-level reasoning and task narration using an integrated LLM. With this addition, the robot could leverage common-sense reasoning to make more intelligent navigation decisions. For example, when prompted to locate an item in a different room, the robot could infer that heading toward a door is the most logical action. Moreover, in bandwidth-limited applications, the LLM could provide real-time task narration by interpreting visual embeddings to describe mission progress. This approach would significantly reduce the communication bandwidth required compared to streaming raw video. By addressing these limitations and pursuing the outlined improvements, the proposed system can evolve into a more robust, efficient, and versatile platform for VLN in complex real-world environments.
7 Conclusion
In this paper, we introduced VL-Explore, a novel navigation pipeline designed for simultaneous exploration and target discovery by autonomous ground robots in unknown environments, leveraging the power of VLMs. Unlike traditional approaches that decouple exploration from path planning and rely on inefficient algorithms, VL-Explore  enables zero-shot inference and efficient navigation using only monocular vision, without prior maps or target-specific information. To validate our approach, we developed a functional prototype UGV platform named Open Rover, optimized for general-purpose VLN tasks in real-world environments. Extensive evaluations demonstrated that VL-Explore  outperforms state-of-the-art map traversal algorithms in efficiency and achieves comparable performance to path-planning methods that rely on prior knowledge. The integration of a CLIP-based VLM into a real-time navigation system underscores the potential of VL-Explore  to advance intelligent robotic exploration. As a modular and flexible framework, VL-Explore  lays the groundwork for future research in applying VLMs to more complex and dynamic robotic applications such as autonomous warehousing, security patrolling, and smart home assistance.
Acknowledgments
This work is supported in part by the National Science Foundation (NSF) grant #$2330416$; the Office of Naval Research (ONR) grants #N000142312429, #N000142312363; and the University of Florida (UF) ROSF research grant #$132763$
Appendix
Appendix A Frame Slicing Strategies
@@
To enhance the embedding of positional information within the pipeline, we preprocess the raw camera frames by slicing them into smaller tiles before inputting them into the vision-language model. This slicing strategy is specifically designed to optimize the robot’s ability to make informed navigation decisions based on its visual perceptions. As illustrated in Fig. 1 in the main paper, the slicer divides each frame into six ($2~{}\text{rows}\times 3~{}\text{columns}$) tiles, with the number of horizontal divisions aligned with the robot’s control capabilities. Using smaller tiles instead of the full camera frame reduces visual distractions, thereby improving the accuracy of the correlation results.
@@
Appendix B Text Prompt Generation
@@
We utilize a YAML-based syntax to define the structure of the prompt databases. YAML, a widely adopted and human-readable markup language, has been extended in our implementation to incorporate additional syntactical features for enhanced flexibility and fine-grained control over the generation of text prompt combinations. Two key features introduced are selective integration and in-place expansion.
@@
    In-place expansion facilitates the collective definition of similar short prompts by using a vertical bar syntax (i.e. |) to separate terms. This feature is particularly effective when used in conjunction with selective integration. For example, the prompt “A photo with no {context|texture|information}” is expanded into three distinct prompts. Although not fitting into the hierarchical template structure, these can be defined as top-level negative prompts, aiding in distinguishing meaningful content from irrelevant or meaningless information.